A man strikes one end of a thin rod with a hammer.
Answer #1

Using relation

d =s*t

Where d is the distance of the rod

s is the speed

t is the time

Since distance is same

sa/sr = tr/ta

1/19 = tr/ta

ta = 19tr

Also ta - tr = 0.16s

equating both equatins

ta = 19(ta - 0.16 )

ta = 19ta - 3.04

18ta = 3.04

ta = 0.1689 s

Hence length of the rod = 345*0.1689 = 58.267 m